OpenStudy (anonymous):

5x-y=23 3x+39=39 Solve the following system by eliminating the y variable. (If the system has no solution, enter NONE for each answer. If the system is dependent, enter INFINITELY MANY for each answer.)

OpenStudy (anonymous):

is there a y variable in the second equation?

OpenStudy (anonymous):

sorry 5x-y=23 3x+3y=39

OpenStudy (anonymous):

5x-23=y and y=x-13